# Moduli spaces of twisted equivariant $G$-bundles over a curve

### v1

@@ 1. Introduction
When $c$ is trivial, the category of $\theta$-twisted $\Gamma$-equivariant bundles appears in the study of the moduli space of $G$-Higgs bundles by Donagi–Gaitsgory [12], and in relation with parahoric bundles in the work of Damiolini [10] and Damiolini–Hong [11]. In full generality, $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles appear naturally in the study of the moduli space of principal bundles with non-connected structure group and non-abelian Hodge theory — see our joint work with García-Prada–Gothen–Mundet i Riera [3], and the work of García-Prada–Gothen–Mundet i Riera [13] —, and also in the study of fixed point varieties of finite group actions on the moduli space of $G$-bundles according to our joint work with García-Prada–Basu [2].
@@ 4. Pseudo-equivariant $G$-bundles as pseudo-equivariant $\rho$-pairs
Following the identification of $G$-bundles with swamps in [20, Section 2.4], we identify pseudo-equivariant $G$-bundles with analogous objects. We still assume that $G$ is a connected semisimple complex Lie group in this section. As in Section 3.1, fix the data $X$, $\Gamma$, $\theta$, $c$ and $\iota:G\times_{(\theta,c)}\Gamma\hookrightarrow\mathrm{SL}(n,\mathbb{C})$.
@@ 4.1. Semistable pseudo-equivariant pairs
where $\sigma|_{\eta}\in\mathbb{W}$, and $\oplus f|_{\eta}$ and $\oplus f^{*-1}|_{\eta}\in\mathbb{H}_{\Gamma}$ — here we use the notation $\oplus f:=\bigoplus_{\gamma\in{\Gamma}}f_{\gamma}$. It is straightforward to check that (14) does not depend on $r$. By [20, Proposition 1.5.1.35] $\mu$ is invariant by the $\operatorname{Stab}_{\mathrm{GL}(n,\mathbb{C}(X))}(\mathbb{E}_{\bullet})\leq\mathrm{GL}(n,\mathbb{C}(X))$-action on $\lambda$, therefore (14) only depends on the data $(E_{\bullet},\alpha_{\bullet},\sigma,f)$.
@@
Given a pseudo-equivariant $\rho$-pair $(E,\sigma,f)$, the underlying pair $(E,\sigma)$ defines a $\rho$-swamp $(E,\mathcal{O}_{X},\sigma)$ in the sense of [20]. Moreover, the swamp $(E,\mathcal{O}_{X},\sigma)$ is $\delta$-(semi)stable, in the sense of [20, Section 2.3.2], if and only if $(E,\sigma,f)$ is $(\delta,0)$-(semi)stable.
@@
We follow the argument in [20, Example 1.5.1.18]. Choose a maximal torus $T\subset\mathrm{GL}(n,\mathbb{C}(X))$ with Lie algebra $\mathfrak{t}$ and a Weyl chamber $K\subset X(\mathfrak{t})$, where $X(\mathfrak{t})$ is the vector space of characters of $\mathfrak{t}$. Given a representation $\rho:\mathrm{GL}(n,\mathbb{C}(X))\to\mathrm{GL}(V)$, let $X(V)\subset X(\mathfrak{t})$ be the set of weights of $\rho$. Given a subset $R\subset X(V)$ and $\chi\in R$, set
@@ 4.3. Semistable pseudo-equivariant $G$-bundles as semistable pseudo-equivariant pairs
To prove the other direction, let $(E,h,f)$ be a (semi)stable pseudo-equivariant $G$-bundle. Let $(E_{\bullet},\alpha_{\bullet})$ be a proper weighted filtration of $E$, assume without loss that $\alpha_{\bullet}$ is integral and let $\lambda:\mathbb{G}_{m}(\mathbb{C}(X))\to\mathrm{GL}(n,\mathbb{C}(X))$ be a 1-parameter subgroup determining $(E_{\bullet},\alpha_{\bullet})$. By Proposition 2.5, the underlying $G$-bundle $(E,h)$ is semistable. By [20, Corollary 2.4.4.6], for $\delta\gg 0$ the underlying $\rho_{a,b,c}$-swamp $(E,\mathcal{O}_{X},\sigma)$ is $\delta$-semistable — in the sense of [20, Section 2.3.2], see Remark 4.5 —, hence $\mu(E_{\bullet},\alpha_{\bullet},\sigma)\geq 0$ for $\delta\gg 0$ by [20, Proposition 2.3.6.5]. Moreover $\mu(E_{\bullet},\alpha_{\bullet},f)\geq 0$ by Lemma 4.12, so the first condition in Definition 4.13 holds.
Now assume that $\mu(E_{\bullet},\alpha_{\bullet},\sigma)=\mu(E_{\bullet},\alpha_{\bullet},f)=0$, i.e. $\mu(\sigma|_{\eta},\lambda)=\mu(f|_{\eta},\lambda)=0$. Then $h|_{\eta}$ is invariant under the $\mathbb{C}^{*}$-action given by $\lambda$. Therefore, $\lambda$ determines a parabolic subgroup $P$ of $G$ and a reduction of structure group $\tau\in H^{0}(X,Q_{G}/P)$ such that the reduction of structure group of $Q=\mathcal{I}so(\mathcal{O}_{X}^{\oplus n},E)$ determined by $E_{\bullet}$ is an extension of structure group of $Q_{\tau}$. By [20, Proposition 2.4.9.1] we may also choose $s\in\mathfrak{g}$ such that $P=P_{s}$ and the 1-parameter subgroup of $\mathrm{GL}(n,\mathbb{C}(X))$ generated by the exponential of $s\otimes 1$ determines the weighted filtration $(E_{\bullet},r\alpha_{\bullet})$ for some $r\in\mathbb{Q}_{\geq 0}$, so that $\deg Q_{G}(\tau,s)=rM(E_{\bullet},\alpha_{\bullet})$. Note that $s\neq 0$, since $E_{\bullet}$ is a proper filtration of $E$. Therefore, since $G$ is semisimple, $P_{s}\subset G$ is a proper parabolic subgroup.
@@
Assume that $(E,\sigma,f)$ is asymptotically (semi)stable. By Lemma 4.14, the corresponding pseudo-equivariant $G$-bundle is (semi)stable, and so is the underlying $G$-bundle by Proposition 2.5. By [20, Corollary 2.4.4.6], for $\delta\gg 0$ the corresponding $\rho_{a,b,c}$-swamp $(E,\mathcal{O}_{X},\sigma)$ is $\delta$-semistable. By [20, Theorem 2.3.4.1] and Remark 3.2, there exists a constant $C$, independent of $(E,\sigma,f)$, such that
@@
By Lemma 4.6, $(\delta,\chi)$-(semi)stability only has to be verified on weighted filtrations $(E_{\bullet},\alpha_{\bullet})$ such that $(\operatorname{rk}E_{\bullet},\alpha_{\bullet})\in S$, for some finite set $S$. We may assume that $S$ only features integral weights. Using (25), we may find a negative number $-\epsilon_{0}$ such that $M(E_{\bullet},\alpha_{\bullet})>-\epsilon_{0}$ for such weighted filtrations. Choose $\delta\geq\epsilon_{0}$. Using [20, Lemma 1.5.1.41], we may find a negative number $-\epsilon_{1}$ such that $M(E_{\bullet},\alpha_{\bullet})+\delta\mu(E_{\bullet},\alpha_{\bullet},\sigma)>-\epsilon_{1}$, for every weighted filtration $(E_{\bullet},\alpha_{\bullet})$ such that $(\operatorname{rk}E_{\bullet},\alpha_{\bullet})\in S$. Choose $\chi\geq\epsilon_{1}$. Then, for such weighted filtrations, since the weights are integral we find
@@ 5.1. Bounded families of vector bundles and the Quot Scheme
Assume that $\chi\gg\delta\gg 0$, so that we may use Proposition 4.11. By [4, Proposition 4.1] the family of semistable twisted equivariant $G$-bundles over $X$ has an underlying family of semistable $G$-bundles, whose associated family $\mathfrak{G}$ of vector bundles of rank $n$ and trivial determinant — obtained via extension of structure group to $\mathrm{GL}(n,\mathbb{C})$ — is bounded by [20, Corollary 2.4.4.6, Proposition 2.2.3.7 and Theorem 2.3.4.1]. Let $\mathcal{O}_{X}(1)$ be a $\Gamma$-invariant ample line bundle on $X$; this can be achieved by taking any ample line bundle and tensoring it by its pullbacks with respect to the elements of $\Gamma$. Set $E(m):=E\otimes\mathcal{O}_{X}(1)^{\otimes m}$, where $E$ is a vector space or a vector bundle over $X$. By [20, Corollary 2.2.3.4] and Remark 3.2, we know that there exist a complex vector space $V$ and a natural number $m$ satisfying the following.
@@ 5.2. The parameter space of $G$-bundles
In [20] the moduli space of holomorphic principal $G$-bundles is constructed using $\rho$-swamps, which are triples $(E,L,\varphi)$ made of a vector bundle $E$, a line bundle $L$ and a surjection $\varphi:E(W)\to L$, where $\rho:\mathrm{GL}(n,\mathbb{C})\to W$ is some representation and $E(W)=\mathcal{H}om(\mathcal{O}_{X}^{\oplus n},E)\times_{\rho}W$ is the associated vector bundle. Indeed, if we ignore the pseudo-equivariant structure in Section 4.2, we obtain an embedding of the category of $G$-bundles into the category of $\rho_{a,b,c}$-pairs $(E,\sigma)$ consisting of a vector bundle $E$ of rank $n$ and degree 00, and a homomorphism $\sigma:E(\mathbb{C}^{n}_{a,b,c})\to\mathcal{O}_{X}$ — these are just $\Gamma$-pseudo-equivariant $\rho_{a,b,c}$-pairs for $\Gamma=1$. In what follows we build a parameter space for this pairs, following [20, Section 2.3].
@@
By [20, Proposition 2.3.5.1], the set of points $t\in\mathfrak{B}$ such that the restriction of $\varphi^{\prime}|_{\{t\}\times X}$ to the kernel of $q_{a,b,c}$ is zero is a closed subscheme ${\mathfrak{I}}_{1}\subset\mathfrak{B}$. On this subscheme, $\varphi^{\prime}$ induces a homomorphism
@@ 5.3. The parameter space of twisted equivariant pairs
are zero on $\{t\}\times X$, for each $\gamma\in\Gamma$. By [20, Proposition 2.3.5.1] this locus is closed, so $\mathfrak{C}_{1}$ is open in $\mathfrak{C}$.
@@
of vector bundles over ×${\mathfrak{I}}\times X$ given by (the pullback of) restricting each homomorphism $V_{\mathfrak{I}}(-m)\to(\operatorname{id}_{\mathfrak{I}}\times\gamma)^{*}V_{\mathfrak{I}}(-m)$ to $\ker q_{\mathfrak{I}}$ and composing with $(\operatorname{id}_{\mathfrak{I}}\times\gamma)^{*}q_{\mathfrak{I}}$. The locus of $\mathfrak{C}$ where $\psi_{0}$ is zero is a closed subscheme $\mathfrak{C}_{2}$ of $\mathfrak{C}$ by [20, Proposition 2.3.5.1]. This is precisely the subvariety where $\psi^{\prime}$ descends to a section $\psi$ of
@@
Again by [20, Proposition 2.3.5.1], the locus of $\mathfrak{C}$ where the image of $\psi^{\prime}$ by the homomorphism
@@
is 00 — note that we are using the notation $\psi^{\prime}=\bigoplus_{\gamma\in\Gamma}\psi^{\prime}_{\gamma}$. By [20, Proposition 2.3.5.1], this is a closed subscheme of $\mathfrak{C}_{1}$, so it is locally closed in $\mathfrak{C}$.
@@ 6.1. Gieseker space and linearization
In order to find an appropriate linearization of the $\mathrm{GL}(V)$-action on $\mathfrak{C}_{0}$ and to apply the Hilbert–Mumford criterion, we use an injective morphism into a projective space, usually called a Gieseker space — the role of Gieseker spaces in the construction of the moduli space of $G$-bundles is explained in [20, Section 2.3.5].
@@ 6.2. Hilbert–Mumford criterion vs (semi)stability
Fix a vector bundle $E\in\mathfrak{G}$ with corresponding point $t\in\mathfrak{C}_{0}$. Following the proof of [20, 2.3.5.12], we may find a lower bound for
@@
Since $F$ is $\Gamma$-invariant, so is $V_{1}$, the weighted filtration $V_{\bullet}(\lambda)$ and $\lambda$ itself. Hence, if $\operatorname{Gies}(t)=(t_{1},t_{2},t_{3})$, then $\mu(t_{3},\lambda)=0$. In particular, $\mu(\operatorname{Gies}(t),\lambda)=s_{1}\mu(t_{1},\lambda)+s_{2}\mu(t_{2},\lambda)$. Since $\operatorname{Gies}(t)$ is semistable, $s_{1}\mu(t_{1},\lambda)+s_{2}\mu(t_{2},\lambda)\geq 0$. Applying the same calculation as in the proof of [20, 2.3.5.12], we conclude that $\mu_{\text{min}}^{\Gamma}(E)\geq-\frac{\delta a}{n}-g$, as required. ∎
@@
for some $i$, then $M(E_{\bullet},\alpha_{\bullet})>C^{\prime}$. This follows directly from $\mu(E_{j})\leq\mu_{\text{max}}^{\Gamma}(E)$ and the uniform bound on $\mu_{\text{max}}^{\Gamma}(E)$ given by Lemma 6.1, where $\mu_{\text{max}}^{\Gamma}(E)$ is the maximum slope attained by a $\Gamma$-invariant vector subbundle of $E$. Using [20, Lemma 1.5.1.41], we may determine a constant $C^{\prime\prime\prime}$ such that $\mu(E_{\bullet},\alpha_{\bullet},\sigma)\geq-C^{\prime\prime\prime}$ for every $\Gamma$-invariant weighted filtration $(E_{\bullet},\alpha_{\bullet})$ such that $(\operatorname{rk}E_{\bullet},\alpha_{\bullet})\in S$. If we choose $C^{\prime}\geq\delta C^{\prime\prime\prime}$, then
@@
Therefore, $(\delta,\chi)$-(semi)stability only has to be verified for $\Gamma$-invariant weighted filtrations $(E_{\bullet},\alpha_{\bullet})$ such that $(\operatorname{rk}E_{\bullet},\alpha_{\bullet})\in S$ and (29) fails. By a slightly refined version of [20, Proposition 2.2.3.7], these weighted filtrations live in bounded families, hence the lemma follows by [20, Proposition 2.2.3.2]. ∎
@@
We just need to check condition (2) in the statement of Lemma 6.2, applied to $(E_{t},\sigma_{t},f_{t})$. To do so, follow the same argument as in the proof of [20, Theorem 2.3.5.13] after the statement of [20, Corollary 2.3.5.14]. ∎
@@
By the proof of [20, Proposition 2.3.5.15], the lemma follows if we can prove that $M(E_{\bullet}^{A},\alpha_{\bullet}^{A})+\delta\mu(E_{\bullet}^{A},\alpha_{\bullet}^{A},\sigma)\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0$ for an arbitrary $\Gamma$-invariant weighted filtration $(E_{\bullet},\alpha_{\bullet})$ — indeed, the last inequality of the last computation in the proof is the only place where the (semi)stability assumptions are used. Note that $\Gamma$-invariance of $(E_{\bullet},\alpha_{\bullet})$ implies, in particular, that $(E_{\bullet}^{A},\alpha_{\bullet}^{A})$ is $\Gamma$-invariant. By Lemma 4.12 and Definition 4.4, the $(\delta,\chi)$-(semi)stability of $(E,\sigma,f)$ implies that $M(E_{\bullet}^{A},\alpha_{\bullet}^{A})+\delta\mu(E_{\bullet}^{A},\alpha_{\bullet}^{A},\sigma)\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0$, as required. ∎
@@
Since $\chi\gg\epsilon^{-1}\gg 0$, it is enough to check that $s_{1}\mu(t_{1},\lambda)+s_{2}\mu(t_{2}),\lambda)\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0$ for all non-trivial $\Gamma$-invariant 1-parameter subgroups $\lambda$. To do this, just copy the proof of [20, Theorem 2.3.5.16] and replace the arbitrary $\lambda$ with a $\Gamma$-invariant $\lambda$, and [20, Proposition 2.3.5.15] by Lemma 6.4. This works because the $\Gamma$-invariance of $\lambda$ clearly implies that the associated filtration $E_{\bullet}$ is $\Gamma$-invariant. ∎

### v2

@@ 1. Introduction
When $c$ is trivial, the moduli stack of $\theta$-twisted $\Gamma$-equivariant $G$-bundles appears in relation with parahoric bundles in the work of Damiolini [10] and Damiolini–Hong [11]. Pappas–Rapoport [20] consider a more general setup where $G$ is replaced by a — non-necessarily constant — group scheme $\mathcal{G}^{\prime}$ over $X$, in relation to tamely ramified bundles. The category of $\theta$-twisted $\Gamma$-equivariant bundles also plays an important role in the study of the moduli space of $G$-Higgs bundles by Donagi–Gaitsgory [12].
In full generality, $(\theta,c)$-twisted $\Gamma$-equivariant $G$-bundles appear naturally in the study of the moduli space of principal bundles with non-connected structure group and non-abelian Hodge theory — see our joint work with García-Prada–Gothen–Mundet i Riera [3], and the work of García-Prada–Gothen–Mundet i Riera [13] —, and also in the study of fixed point varieties of finite group actions on the moduli space of $G$-bundles according to our joint work with García-Prada–Basu [2].
@@ 4. Pseudo-equivariant $G$-bundles as pseudo-equivariant $\rho$-pairs
Following the identification of $G$-bundles with swamps in [21, Section 2.4], we identify pseudo-equivariant $G$-bundles with analogous objects. We still assume that $G$ is a connected semisimple complex Lie group in this section. As in Section 3.1, fix the data $X$, $\Gamma$, $\theta$, $c$ and $\iota:G\times_{(\theta,c)}\Gamma\hookrightarrow\mathrm{SL}(n,\mathbb{C})$.
@@ 4.1. Semistable pseudo-equivariant pairs
where $\sigma|_{\eta}\in\mathbb{W}$, and $\oplus f|_{\eta}$ and $\oplus f^{*-1}|_{\eta}\in\mathbb{H}_{\Gamma}$ — here we use the notation $\oplus f:=\bigoplus_{\gamma\in{\Gamma}}f_{\gamma}$. It is straightforward to check that (14) does not depend on $r$. By [21, Proposition 1.5.1.35] $\mu$ is invariant by the $\operatorname{Stab}_{\mathrm{GL}(n,\mathbb{C}(X))}(\mathbb{E}_{\bullet})\leq\mathrm{GL}(n,\mathbb{C}(X))$-action on $\lambda$, therefore (14) only depends on the data $(E_{\bullet},\alpha_{\bullet},\sigma,f)$.
@@
Given a pseudo-equivariant $\rho$-pair $(E,\sigma,f)$, the underlying pair $(E,\sigma)$ defines a $\rho$-swamp $(E,\mathcal{O}_{X},\sigma)$ in the sense of [21]. Moreover, the swamp $(E,\mathcal{O}_{X},\sigma)$ is $\delta$-(semi)stable, in the sense of [21, Section 2.3.2], if and only if $(E,\sigma,f)$ is $(\delta,0)$-(semi)stable.
@@
We follow the argument in [21, Example 1.5.1.18]. Choose a maximal torus $T\subset\mathrm{GL}(n,\mathbb{C}(X))$ with Lie algebra $\mathfrak{t}$ and a Weyl chamber $K\subset X(\mathfrak{t})$, where $X(\mathfrak{t})$ is the vector space of characters of $\mathfrak{t}$. Given a representation $\rho:\mathrm{GL}(n,\mathbb{C}(X))\to\mathrm{GL}(V)$, let $X(V)\subset X(\mathfrak{t})$ be the set of weights of $\rho$. Given a subset $R\subset X(V)$ and $\chi\in R$, set
@@ 4.3. Semistable pseudo-equivariant $G$-bundles as semistable pseudo-equivariant pairs
To prove the other direction, let $(E,h,f)$ be a (semi)stable pseudo-equivariant $G$-bundle. Let $(E_{\bullet},\alpha_{\bullet})$ be a proper weighted filtration of $E$, assume without loss that $\alpha_{\bullet}$ is integral and let $\lambda:\mathbb{G}_{m}(\mathbb{C}(X))\to\mathrm{GL}(n,\mathbb{C}(X))$ be a 1-parameter subgroup determining $(E_{\bullet},\alpha_{\bullet})$. By Proposition 2.5, the underlying $G$-bundle $(E,h)$ is semistable. By [21, Corollary 2.4.4.6], for $\delta\gg 0$ the underlying $\rho_{a,b,c}$-swamp $(E,\mathcal{O}_{X},\sigma)$ is $\delta$-semistable — in the sense of [21, Section 2.3.2], see Remark 4.5 —, hence $\mu(E_{\bullet},\alpha_{\bullet},\sigma)\geq 0$ for $\delta\gg 0$ by [21, Proposition 2.3.6.5]. Moreover $\mu(E_{\bullet},\alpha_{\bullet},f)\geq 0$ by Lemma 4.12, so the first condition in Definition 4.13 holds.
Now assume that $\mu(E_{\bullet},\alpha_{\bullet},\sigma)=\mu(E_{\bullet},\alpha_{\bullet},f)=0$, i.e. $\mu(\sigma|_{\eta},\lambda)=\mu(f|_{\eta},\lambda)=0$. Then $h|_{\eta}$ is invariant under the $\mathbb{C}^{*}$-action given by $\lambda$. Therefore, $\lambda$ determines a parabolic subgroup $P$ of $G$ and a reduction of structure group $\tau\in H^{0}(X,Q_{G}/P)$ such that the reduction of structure group of $Q=\mathcal{I}so(\mathcal{O}_{X}^{\oplus n},E)$ determined by $E_{\bullet}$ is an extension of structure group of $Q_{\tau}$. By [21, Proposition 2.4.9.1] we may also choose $s\in\mathfrak{g}$ such that $P=P_{s}$ and the 1-parameter subgroup of $\mathrm{GL}(n,\mathbb{C}(X))$ generated by the exponential of $s\otimes 1$ determines the weighted filtration $(E_{\bullet},r\alpha_{\bullet})$ for some $r\in\mathbb{Q}_{\geq 0}$, so that $\deg Q_{G}(\tau,s)=rM(E_{\bullet},\alpha_{\bullet})$. Note that $s\neq 0$, since $E_{\bullet}$ is a proper filtration of $E$. Therefore, since $G$ is semisimple, $P_{s}\subset G$ is a proper parabolic subgroup.
@@
Assume that $(E,\sigma,f)$ is asymptotically (semi)stable. By Lemma 4.14, the corresponding pseudo-equivariant $G$-bundle is (semi)stable, and so is the underlying $G$-bundle by Proposition 2.5. By [21, Corollary 2.4.4.6], for $\delta\gg 0$ the corresponding $\rho_{a,b,c}$-swamp $(E,\mathcal{O}_{X},\sigma)$ is $\delta$-semistable. By [21, Theorem 2.3.4.1] and Remark 3.2, there exists a constant $C$, independent of $(E,\sigma,f)$, such that
@@
By Lemma 4.6, $(\delta,\chi)$-(semi)stability only has to be verified on weighted filtrations $(E_{\bullet},\alpha_{\bullet})$ such that $(\operatorname{rk}E_{\bullet},\alpha_{\bullet})\in S$, for some finite set $S$. We may assume that $S$ only features integral weights. Using (25), we may find a negative number $-\epsilon_{0}$ such that $M(E_{\bullet},\alpha_{\bullet})>-\epsilon_{0}$ for such weighted filtrations. Choose $\delta\geq\epsilon_{0}$. Using [21, Lemma 1.5.1.41], we may find a negative number $-\epsilon_{1}$ such that $M(E_{\bullet},\alpha_{\bullet})+\delta\mu(E_{\bullet},\alpha_{\bullet},\sigma)>-\epsilon_{1}$, for every weighted filtration $(E_{\bullet},\alpha_{\bullet})$ such that $(\operatorname{rk}E_{\bullet},\alpha_{\bullet})\in S$. Choose $\chi\geq\epsilon_{1}$. Then, for such weighted filtrations, since the weights are integral we find
@@ 5.1. Bounded families of vector bundles and the Quot Scheme
Assume that $\chi\gg\delta\gg 0$, so that we may use Proposition 4.11. By [4, Proposition 4.1] the family of semistable twisted equivariant $G$-bundles over $X$ has an underlying family of semistable $G$-bundles, whose associated family $\mathfrak{G}$ of vector bundles of rank $n$ and trivial determinant — obtained via extension of structure group to $\mathrm{GL}(n,\mathbb{C})$ — is bounded by [21, Corollary 2.4.4.6, Proposition 2.2.3.7 and Theorem 2.3.4.1]. Let $\mathcal{O}_{X}(1)$ be a $\Gamma$-invariant ample line bundle on $X$; this can be achieved by taking any ample line bundle and tensoring it by its pullbacks with respect to the elements of $\Gamma$. Set $E(m):=E\otimes\mathcal{O}_{X}(1)^{\otimes m}$, where $E$ is a vector space or a vector bundle over $X$. By [21, Corollary 2.2.3.4] and Remark 3.2, we know that there exist a complex vector space $V$ and a natural number $m$ satisfying the following.
@@ 5.2. The parameter space of $G$-bundles
In [21] the moduli space of holomorphic principal $G$-bundles is constructed using $\rho$-swamps, which are triples $(E,L,\varphi)$ made of a vector bundle $E$, a line bundle $L$ and a surjection $\varphi:E(W)\to L$, where $\rho:\mathrm{GL}(n,\mathbb{C})\to W$ is some representation and $E(W)=\mathcal{H}om(\mathcal{O}_{X}^{\oplus n},E)\times_{\rho}W$ is the associated vector bundle. Indeed, if we ignore the pseudo-equivariant structure in Section 4.2, we obtain an embedding of the category of $G$-bundles into the category of $\rho_{a,b,c}$-pairs $(E,\sigma)$ consisting of a vector bundle $E$ of rank $n$ and degree 00, and a homomorphism $\sigma:E(\mathbb{C}^{n}_{a,b,c})\to\mathcal{O}_{X}$ — these are just $\Gamma$-pseudo-equivariant $\rho_{a,b,c}$-pairs for $\Gamma=1$. In what follows we build a parameter space for this pairs, following [21, Section 2.3].
@@
By [21, Proposition 2.3.5.1], the set of points $t\in\mathfrak{B}$ such that the restriction of $\varphi^{\prime}|_{\{t\}\times X}$ to the kernel of $q_{a,b,c}$ is zero is a closed subscheme ${\mathfrak{I}}_{1}\subset\mathfrak{B}$. On this subscheme, $\varphi^{\prime}$ induces a homomorphism
@@ 5.3. The parameter space of twisted equivariant pairs
are zero on $\{t\}\times X$, for each $\gamma\in\Gamma$. By [21, Proposition 2.3.5.1] this locus is closed, so $\mathfrak{C}_{1}$ is open in $\mathfrak{C}$.
@@
of vector bundles over ×${\mathfrak{I}}\times X$ given by (the pullback of) restricting each homomorphism $V_{\mathfrak{I}}(-m)\to(\operatorname{id}_{\mathfrak{I}}\times\gamma)^{*}V_{\mathfrak{I}}(-m)$ to $\ker q_{\mathfrak{I}}$ and composing with $(\operatorname{id}_{\mathfrak{I}}\times\gamma)^{*}q_{\mathfrak{I}}$. The locus of $\mathfrak{C}$ where $\psi_{0}$ is zero is a closed subscheme $\mathfrak{C}_{2}$ of $\mathfrak{C}$ by [21, Proposition 2.3.5.1]. This is precisely the subvariety where $\psi^{\prime}$ descends to a section $\psi$ of
@@
Again by [21, Proposition 2.3.5.1], the locus of $\mathfrak{C}$ where the image of $\psi^{\prime}$ by the homomorphism
@@
is 00 — note that we are using the notation $\psi^{\prime}=\bigoplus_{\gamma\in\Gamma}\psi^{\prime}_{\gamma}$. By [21, Proposition 2.3.5.1], this is a closed subscheme of $\mathfrak{C}_{1}$, so it is locally closed in $\mathfrak{C}$.
@@ 6.1. Gieseker space and linearization
In order to find an appropriate linearization of the $\mathrm{GL}(V)$-action on $\mathfrak{C}_{0}$ and to apply the Hilbert–Mumford criterion, we use an injective morphism into a projective space, usually called a Gieseker space — the role of Gieseker spaces in the construction of the moduli space of $G$-bundles is explained in [21, Section 2.3.5].
@@ 6.2. Hilbert–Mumford criterion vs (semi)stability
Fix a vector bundle $E\in\mathfrak{G}$ with corresponding point $t\in\mathfrak{C}_{0}$. Following the proof of [21, 2.3.5.12], we may find a lower bound for
@@
Since $F$ is $\Gamma$-invariant, so is $V_{1}$, the weighted filtration $V_{\bullet}(\lambda)$ and $\lambda$ itself. Hence, if $\operatorname{Gies}(t)=(t_{1},t_{2},t_{3})$, then $\mu(t_{3},\lambda)=0$. In particular, $\mu(\operatorname{Gies}(t),\lambda)=s_{1}\mu(t_{1},\lambda)+s_{2}\mu(t_{2},\lambda)$. Since $\operatorname{Gies}(t)$ is semistable, $s_{1}\mu(t_{1},\lambda)+s_{2}\mu(t_{2},\lambda)\geq 0$. Applying the same calculation as in the proof of [21, 2.3.5.12], we conclude that $\mu_{\text{min}}^{\Gamma}(E)\geq-\frac{\delta a}{n}-g$, as required. ∎
@@
for some $i$, then $M(E_{\bullet},\alpha_{\bullet})>C^{\prime}$. This follows directly from $\mu(E_{j})\leq\mu_{\text{max}}^{\Gamma}(E)$ and the uniform bound on $\mu_{\text{max}}^{\Gamma}(E)$ given by Lemma 6.1, where $\mu_{\text{max}}^{\Gamma}(E)$ is the maximum slope attained by a $\Gamma$-invariant vector subbundle of $E$. Using [21, Lemma 1.5.1.41], we may determine a constant $C^{\prime\prime\prime}$ such that $\mu(E_{\bullet},\alpha_{\bullet},\sigma)\geq-C^{\prime\prime\prime}$ for every $\Gamma$-invariant weighted filtration $(E_{\bullet},\alpha_{\bullet})$ such that $(\operatorname{rk}E_{\bullet},\alpha_{\bullet})\in S$. If we choose $C^{\prime}\geq\delta C^{\prime\prime\prime}$, then
@@
Therefore, $(\delta,\chi)$-(semi)stability only has to be verified for $\Gamma$-invariant weighted filtrations $(E_{\bullet},\alpha_{\bullet})$ such that $(\operatorname{rk}E_{\bullet},\alpha_{\bullet})\in S$ and (29) fails. By a slightly refined version of [21, Proposition 2.2.3.7], these weighted filtrations live in bounded families, hence the lemma follows by [21, Proposition 2.2.3.2]. ∎
@@
We just need to check condition (2) in the statement of Lemma 6.2, applied to $(E_{t},\sigma_{t},f_{t})$. To do so, follow the same argument as in the proof of [21, Theorem 2.3.5.13] after the statement of [21, Corollary 2.3.5.14]. ∎
@@
By the proof of [21, Proposition 2.3.5.15], the lemma follows if we can prove that $M(E_{\bullet}^{A},\alpha_{\bullet}^{A})+\delta\mu(E_{\bullet}^{A},\alpha_{\bullet}^{A},\sigma)\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0$ for an arbitrary $\Gamma$-invariant weighted filtration $(E_{\bullet},\alpha_{\bullet})$ — indeed, the last inequality of the last computation in the proof is the only place where the (semi)stability assumptions are used. Note that $\Gamma$-invariance of $(E_{\bullet},\alpha_{\bullet})$ implies, in particular, that $(E_{\bullet}^{A},\alpha_{\bullet}^{A})$ is $\Gamma$-invariant. By Lemma 4.12 and Definition 4.4, the $(\delta,\chi)$-(semi)stability of $(E,\sigma,f)$ implies that $M(E_{\bullet}^{A},\alpha_{\bullet}^{A})+\delta\mu(E_{\bullet}^{A},\alpha_{\bullet}^{A},\sigma)\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0$, as required. ∎
@@
Since $\chi\gg\epsilon^{-1}\gg 0$, it is enough to check that $s_{1}\mu(t_{1},\lambda)+s_{2}\mu(t_{2}),\lambda)\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle($}}\geq\mathrel{\raisebox{-2.15277pt}{$\scriptscriptstyle)$}}0$ for all non-trivial $\Gamma$-invariant 1-parameter subgroups $\lambda$. To do this, just copy the proof of [21, Theorem 2.3.5.16] and replace the arbitrary $\lambda$ with a $\Gamma$-invariant $\lambda$, and [21, Proposition 2.3.5.15] by Lemma 6.4. This works because the $\Gamma$-invariance of $\lambda$ clearly implies that the associated filtration $E_{\bullet}$ is $\Gamma$-invariant. ∎